### v2

@@ 3.3 Dynamic Matrices
At last, we outline how to handle matrices $\mathbf{M}$ that are allowed to update over time, i.e., Theorem 1.2. See 1.2
We sketch how to handle column insertions: Let $\mathbf{M}$ be the current matrix. We split it into smaller matrices $\mathbf{M}^{(0)},\mathbf{M}^{(1)},...,\mathbf{M}^{(\log n)}$, where each $\mathbf{M}^{(i)}$ contains at most $2^{i}$ columns of $\mathbf{M}$. When a new column is inserted to $\mathbf{M}$, we insert that column to $\mathbf{M}^{(0)}$. When any $\mathbf{M}^{(i)}$ contains more than $2^{i}$ columns, then all its columns are inserted to $\mathbf{M}^{(i+1)}$ and $\mathbf{M}^{(i)}$ is reset to an empty matrix. Thus, any $\mathbf{M}^{(i)}$ receives new columns at most once every $2^{i}$ iterations to $\mathbf{M}$.
When an $\mathbf{M}^{(i)}$ receives new columns, we initialize a new matrix-vector multiplication data structure (Theorem 1.1) on it. This takes $\widetilde{O}(m2^{i})$ time, but amortizes to $\widetilde{O}(m)$ time per insertion to $\mathbf{M}$ since it only happens at most once every $2^{i}$ insertions. This yields an amortized time of $\sum_{i=0}^{\log n}\widetilde{O}(m)=\widetilde{O}(m)$ time to maintain all $\mathbf{M}^{(i)}$ for $i=0,...,\log n$.
To answer a matrix-vector multiplication query $\mathbf{M}v$, we split $v$ into vectors $v^{(0)},...,v^{(\log n)}$ such that $\mathbf{M}v=\sum_{i=0}^{\log n}\mathbf{M}^{(i)}v^{(i)}$. For the time complexity, note that each $\mathbf{M}^{(i)}$ is a submatrix of $\mathbf{M}$. Thus, the the time complexity needed to multiply a vector by $\mathbf{M}^{(i)}$ is bounded by $\widetilde{O}(mn^{1-1/d})$.
For column deletions, we can ignore them until $\mathbf{M}^{(i)}$ is reinitialized: simply set $v_{j}=0$ for the columns that should have been deleted. We can handle row deletions updates to $\mathbf{M}$ in a similar way, by splitting the matrix again into $\log m$ matrices, each containing at most $2^{j}$ rows. The full proof is given in Section A.2.
@@ 4 Conclusion
In this work, we study the structural complexity of sequential matrix-vector multiplication. We propose a theoretical framework to study heuristic algorithms that have achieved tremendous practical success. Building on the VC-dimension literature, we propose the notion of a corrupted VC-dimension $d$ and provide structural characterizations in section 1. In theorem 1.1, we show that if a matrix $\mathbf{M}\in\{0,1\}^{m\times n}$ has corrupted VC-dimension $d$, then its matrix-vector product $\mathbf{M}v$ can be computed in $O(mn^{1-1/d})$ time, providing polynomial speed-ups over existing methods. Moreover, our algorithm maintains updates to $\mathbf{M}$ in $\widetilde{O}(m)$ and $\widetilde{O}(n)$ time. This leads to the first $O(n^{2-\epsilon})$-time algorithms for maintaining high-accuracy estimates for Laplacian solvers, effective resistance, and triangle detection in graphs with node insertions and deletions.
Our results motivate interesting future work: for instance, it would be interesting if alternate generalizations of the VC-dimension, such as the Natarajan dimension, can also parameterize the structural complexity of matrix-vector multiplication. Moreover, we conjecture that our algorithm can be used to provide combinatorial accelerations to the context-free grammar parsing and the $k$-clique detection problem in structured instances. Finally, for the non-Boolean setting where we provide subquadratic runtimes in matrices with bounded Pollard pseudodimension, we conjecture that the dependence of the number of thresholds $A$ in the runtime can be reduced to be polylogarithmic instead of linear.
5 Acknowledgements
This work was supported by NSF Grants DMS 2202961 and CCF 2338816. We also express our thanks to Ce Jin, Shunhua Jiang, Elia Gorokhovsky, and Jingtong Sun for insightful discussions.
Appendix A Online Matrix Vector Data Structure
In this section we prove all the intermediate lemmas required for our first two main results on subquadratic online matrix-vector multiplication. The proof that combined the technical lemmas was already given in Section 3.
See 1.1
See 1.2
A.1 Static Online Matrix Vector Multiplication Data Structure
@@
We inductively prove that $\Phi=\mathbf{M}v$ over the number of visited vertices. This is true when only one vertex is visited, as that is the root $r$ and by definition $\Phi_{r}=(\mathbf{M}v)_{r}$.
For the inductive step, assume DFS visits the $(k+1)$st vertex $x$. Let $y$ be the parent of the vertex, then the algorithm assigns
@@
Runtime analysis. Let $\delta_{i,j}$ be the number of non-zero entries in $\Delta_{i,j}$ for all $(i,j)\in E(T)$. Initializing the vector $\Phi$ and populating its first element $\Phi_{r}$ takes $O(n+m)$ time. The cost of populating the remaining entries is proportional to the number of non-zero entries in each $\Delta_{x,y}$, since we can store each $\Delta$ via a sparser representation of only its non-zero entries and use this to compute $\Delta_{y,x}^{\top}v$ in time $O(\delta_{y,x})$. Thus, the total time complexity is $O(\sum_{(x,y)\in E(T)}\delta_{x,y})=O(\mathrm{weight}(T))$. Hence, the runtime of algorithm 1 is $O(\mathrm{weight}(T)+n+m)$, proving the lemma.
@@
Definition A.2 (Path-$\Delta$ matrix $\mathbf{N}^{(T)}$).
@@
which proves the claim.
Lemma A.5.
@@
Suppose the statement holds for matrices of width $<n$ and we now have a matrix $\mathbf{M}$ of width $n$. Let $L\subset V$ be the leaves of $T$. Let $T^{\prime}$ be the tree $T$ with these leafs removed, and let $\mathbf{M}^{\prime}\in\{0,1\}^{m\times(V\setminus L)}$ be matrix $\mathbf{M}$ with all columns corresponding to $L$ removed. Then for any $y\in V\setminus L$
@@
where equation 1 uses the fact that the path from root $r$ to $y$ is a path from root $r$ to $z$ with one extra edge $(z,y)$. Equation 2 uses the fact the path from $r$ to $z$ in $T$ also exists in $T^{\prime}$, and finally equation 3 follows by the induction hypothesis. Together, they prove the lemma.
@@
We compute $\mathbf{N}^{(T)}v$ via Lemma A.3 which requires us to compute the vector $\vec{\sigma}\in\mathbb{R}^{E}$ with $\vec{\sigma}_{(y,c)}=\sum_{z\in V(T_{c})}v_{z}$ for all $(y,c)\in E$ where $T_{c}$ is the subtree of $T$ rooted at $c$. This can be done in $O(n)$ time via a simple “dynamic programming on trees” algorithm: For any $c\in V$, let $C$ be its children. Then
@@
Correctness.
Without loss of generality, assume that for root $r$ of tree $T$ we have $\mathbf{M}_{r}=\vec{0}$. Otherwise, append a $\vec{0}$-column to $\mathbf{M}$, and add edge $(n+1,r)$ to $T$ with label $\Delta_{n+1,r}=\mathbf{M}_{r}-\mathbf{M}_{n+1}=\mathbf{M}_{r}$, and make $n+1$ the new root of $T$. This increases the tree weight by at most $m$, resulting in an additive $O(m)$ in the time complexity.
Append another entry to $v$ so that the dimensions match and $\mathbf{M}v$ is well-defined. By lemma A.5, we thus have $\mathbf{N}^{(T)}v=\mathbf{M}v$.
A.2 Dynamic OMv Data Structure
In this section we give a black-box reduction from dynamic OMv to static OMv. In the static version, we are given a matrix $\mathbf{M}$ to preprocess. This matrix stays fixed, while we receive an online sequence of vectors and must repeatedly compute $\mathbf{M}v$. In the dynamic version, we can inform the data structure about changes to $\mathbf{M}$. In particular, we are allowed to add/remove/change rows and columns of the matrix.
Using the reduction from dynamic to static, we obtain Theorem 1.2.
See 1.2
@@ A.2 Dynamic OMv Data Structure
When there have been $22^{i}/$ 2 postponed deletions to any $\mathbf{M}^{(i)}$, then the columns are actually deleted and the respective static OMv data structure is reinitialized on the new $\mathbf{M}^{(i)}$. The remaining columns in $\mathbf{M}^{(i)}$ all exist in the current dynamic input $\mathbf{M}$, so again, the static OMv data structure receives a submatrix of $\mathbf{M}$ as input. For any one $i$, this takes amortized $O(P(n,2^{i})/2^{i})$ time which is subsumed by the complexity of handling insertions.
Observe that internally, our matrices $\mathbf{M}^{(i)}$ are always at most twice their intended size because of the yet to be removed columns. Thus the query complexity increases by at most a constant factor.
Proof A.11.
of Lemma A.8 (Row & Column Updates)   Here, we extend Lemma A.9 to support row updates as well. The reduction follows the same idea as Lemma A.9.
Let $\mathbf{N}$ be the dynamic input matrix. We run $\log n$ copies of the column update data structure Lemma A.9 on matrices $\mathbf{N}^{(0)},...,\mathbf{N}^{(\log n)}$, with the $i$th matrix having at most $2^{i}$ rows. When inserting new rows to $\mathbf{N}$, they are passed to $\mathbf{N}^{(0)}$. When $\mathbf{N}^{(i)}$ has more than $2^{i}$ rows, all rows are appended to $\mathbf{N}^{(i+1)}$. When rows are appended to any $\mathbf{N}^{(i)}$, the respective data structure is reinitialized. To handle deletions, we remove the respective row from the output until $22^{i}/$ 2 rows have been deleted from $\mathbf{N}^{(i)}$, then the rows are removed from $\mathbf{N}^{(i)}$ and the respective data structure is reinitialized.
The amortized row update time thus becomes
@@
$O\left(\sum_{j=0}^{\log m}\sum_{i=0}^{\log n}P(2^{j},2^{i})/2^{i}\right)=O\left(\sum_{i=0}^{\log n}P(m,2^{i})/2^{i}\right)$
@@
which proves the lemma.
Proof A.12.
of Theorem 1.2 (Dynamic OMv for corrupted VC-dimension $d$)   The initialization cost of static OMv is $P(m,n)=\widetilde{O}(mn)$, and the query complexity is $Q(m,n)=\widetilde{O}(mn^{1-1/d})$ by Theorem 1.1. This complexity also holds for submatrices, i.e., $Q(a,b)=\widetilde{O}(mn^{2-1/d}+n)$ for any $a\times b$ sized submatrix.
@@
$O\left(\sum_{i=0}^{\log n}P(m,2^{i})/2^{i}\right)=\widetilde{O}\left(\sum_{i=0}^{\log n}m2^{i}/2^{i}\right)=\widetilde{O}(m)$
@@
$O\left(\sum_{j=0}^{\log m}P(2^{j},n)/2^{j}\right)=\widetilde{O}\left(\sum_{j=0}^{\log m}2^{j}n/2^{j}\right)=\widetilde{O}(n)$
@@
$\widetilde{O}(Q(m,n))\leq\widetilde{O}(mn^{1-1/d}+n),$
which proves the theorem.
A.3 Extension to Structured Non-Boolean Matrices
In this section, we prove the following lemma.
Lemma A.13.
Suppose that the Pollard pseudodimension of the matrix $\mathbf{M}\in\mathbb{R}^{m\times n}$ is $d$, and that the number of thresholds is $A$. Then, after an $\widetilde{O}(Amn)$ time preprocessing, there is a data structure that, upon receiving a vector $v\in\mathbb{R}^{n}$, can return $\mathbf{M}v$ in time $O(Amn^{1-1/d}+An)$.

### v3

@@ 3.3 Dynamic Matrices
At last, we outline how to handle matrices $\mathbf{M}$ that receive updates over time, i.e., Theorem˜1.4. See 1.4
We sketch how to handle column insertions: Let $\mathbf{M}$ be the current matrix. We split it into smaller matrices $\mathbf{M}^{(0)},\mathbf{M}^{(1)},...,\mathbf{M}^{(\log n)}$, where each $\mathbf{M}^{(i)}$ contains at most $2^{i}$ columns of $\mathbf{M}$. When a new column is inserted to $\mathbf{M}$, we insert that column to $\mathbf{M}^{(0)}$. When any $\mathbf{M}^{(i)}$ contains more than $2^{i}$ columns, then all its columns are inserted to $\mathbf{M}^{(i+1)}$ and $\mathbf{M}^{(i)}$ is reset to an empty matrix. Thus, any $\mathbf{M}^{(i)}$ receives new columns at most once every $2^{i}$ iterations to $\mathbf{M}$. When an $\mathbf{M}^{(i)}$ receives new columns, we initialize a new matrix-vector multiplication data structure (Theorem˜1.3) on it. This takes $\widetilde{O}(m2^{i})$ time, but amortizes to $\widetilde{O}(m)$ time per insertion to $\mathbf{M}$ since it only happens at most once every $2^{i}$ insertions. This yields an amortized time of $\sum_{i=0}^{\log n}\widetilde{O}(m)=\widetilde{O}(m)$ time to maintain all $\mathbf{M}^{(i)}$ for $i=0,...,\log n$. To answer a matrix-vector multiplication query $\mathbf{M}v$, we split $v$ into vectors $v^{(0)},...,v^{(\log n)}$ such that $\mathbf{M}v=\sum_{i=0}^{\log n}\mathbf{M}^{(i)}v^{(i)}$. For the time complexity, note that each $\mathbf{M}^{(i)}$ is a submatrix of $\mathbf{M}$. Thus, the the time complexity needed to multiply a vector by $\mathbf{M}^{(i)}$ is bounded by $\widetilde{O}(nm^{1-1/d}+m)$.
We ignore column deletions to $\mathbf{M}^{(i)}$ until $\mathbf{M}^{(i)}$ is reinitialized: simply set $v_{j}=0$ for the columns that should have been deleted. We can handle row updates to $\mathbf{M}$ in a similar way, by splitting the matrix again into $\log m$ matrices, each containing at most $2^{j}$ rows. The full proof is given in Section˜B.3.
@@ 4 Conclusion
We study the structural complexity of matrix-vector multiplication. We propose a theoretical framework to study heuristic algorithms that have achieved tremendous practical success. Building on the VC-dimension literature, we propose the notion of a corrupted VC-dimension $d$ and provide structural characterizations. We show that if a matrix $\mathbf{M}\in\{0,1\}^{m\times n}$ has corrupted VC-dimension $d$, then its matrix-vector product $\mathbf{M}v$ can be computed in $\widetilde{O}(nm^{1-1/d}+m)$ time, providing polynomial speed-ups over existing methods. Moreover, our algorithm maintains updates to $\mathbf{M}$ in $\widetilde{O}(m)$ and $\widetilde{O}(n)$ time, leading to the first $\widetilde{O}(n^{2-\epsilon})$-time algorithms for maintaining high-accuracy estimates for Laplacian solvers, effective resistance, and triangle detection in dynamic graphs.
Limitations. While we show subquadratic runtimes in non-Boolean matrices with bounded Pollard pseudodimension, we believe that the linear dependence of the number of distinct values $T$ in the runtime is suboptimal and should instead be logarithmic. Finally, without matching lower bounds, we cannot verify whether the $\widetilde{O}(n^{2-1/d})$ scaling with respect to the VC-dimension is optimal.
Societal Impacts. This work is foundational in nature. As such, while it enables more efficient computation in iterative algorithms, it is not tied to any specific applications or deployments.
Acknowledgments and Disclosure of Funding
This work was supported by NSF Grants DMS 2202961 and CCF 2338816. We also express our thanks to Ce Jin, Andrzej Lingas, Albert Weng, Jesper Jansson, Shunhua Jiang, Xiaobai Sun, Elia Gorokhovsky, and Jingtong Sun for insightful discussions.
Appendix A Preliminaries for Supplementary Material
[…]
VC-Dimension of Matrices. Any matrix $\mathbf{M}\in\{0,1\}^{m\times n}$ corresponds to a set family given by $\mathcal{F}_{\mathbf{M}}=\{\{j:\mathbf{M}_{i,j}=1\}:i\in[m]\}$, i.e., consider each row of $\mathbf{M}$ as the indicator vector of some set. Through it, we can encode $\mathbf{M}$ by the range space $\mathfrak{R}_{M}=([n],\mathcal{F}_{\mathbf{M}})$. Then, the VC-dimension of $\mathbf{M}$ is defined as the VC-dimension of $\mathfrak{R}_{\mathbf{M}}$. Here, a subset $S$ of the columns of $\mathbf{M}$ is said to be shattered if each of the $2^{|S|}$ many 0/1 strings appears in some row in the restriction of $\mathbf{M}$ to the columns in $S$. Then, the VC-dimension of $\mathbf{M}$ is the maximum size of a shattered subset of the columns of $\mathbf{M}$. Moreover, the VC-dimension $d_{m,n}$ of a matrix $\mathbf{M}\in\{0,1\}^{m\times n}$ satisfies $1\leq d_{m,n}\leq\log m$.
Pollard Pseudodimension of Matrices. For a family of functions $\mathcal{F}$ mapping $X$ to $Y$, the Pollard pseudodimension of $\mathcal{F}$, denoted by $\mathrm{Pdim}(\mathcal{F})$, is the VC dimension of the set system $(X\times Y,\{\{(x,y)\mid f(x)\geq y\}\mid f\in\mathcal{F}\})$.
For a real matrix $\mathbf{M}$, define the pseudo-dimension of $\mathbf{M}$, denoted by $\mathrm{Pdim}(\mathbf{M})$ by thinking of the rows of $\mathbf{M}$ as functions and taking the pseudo-dimension of the resulting class of functions. Specifically, if $\mathbf{M}$ is an $m\times n$ matrix, for each $i\in\{1,\dots,m\}$, define $f_{\mathbf{M},i}:\{1,\dots,n\}\to\mathbb{R}$ by $f_{\mathbf{M},i}(j)=\mathbf{M}_{i,j}$, and let $\mathrm{Pdim}(\mathbf{M})=\mathrm{Pdim}(\{f_{\mathbf{M},i}:i\in\{1,\dots,m\}\})$.
Appendix B Online Matrix Vector Data Structure
In this section we prove all the intermediate lemmas required for our first two main results on subquadratic online matrix-vector multiplication. The proof that combined the technical lemmas was already given in Section˜3.
See 1.3
See 1.4
B.1 Static Online Matrix Vector Multiplication Data Structure
@@
We inductively prove that $\Phi=\mathbf{M}v$ over the number of visited vertices. This is true when only one vertex is visited, as that is the root $r$ and by definition $\Phi_{r}=(\mathbf{M}v)_{r}$. For the inductive step, assume DFS visits the (k+1)(k+1)st vertex $x$. Let $y$ be the parent of the vertex, then the algorithm assigns
@@
Runtime analysis. Let $\delta_{i,j}$ be the number of non-zero entries in $\Delta_{i,j}$ for all $(i,j)\in E(T)$. Initializing the vector $\Phi$ and populating its first element $\Phi_{r}$ takes $O(n+m)$ time. The cost of populating the remaining entries is proportional to the number of non-zero entries in each $\Delta_{x,y}$, since we can store each $\Delta$ via a sparser representation of only its non-zero entries and use this to compute $\Delta_{y,x}^{\top}v$ in time $O(\delta_{y,x})$. Thus, the total time complexity is $O(\sum_{(x,y)\in E(T)}\delta_{x,y})=O(\mathrm{weight}(T))$. Hence, the runtime of Algorithm˜1 is $O(\mathrm{weight}(T)+n+m)$, proving the lemma. ∎
@@
When the spanning tree is constructed over the columns of $\mathbf{M}$, we run a different algorithm. The following matrix $\mathbf{N}$ is defined based on the spanning tree. We will prove (i) that multiplying vectors with $\mathbf{N}$ can be done efficiently (Lemma˜B.2), and (ii) that $\mathbf{N}=\mathbf{M}$ (Lemma˜B.3). Combining the two then yields Lemma˜3.4.
Definition B.1 (Path-$\Delta$ matrix $\mathbf{N}^{(T)}$).
@@
which proves the claim. ∎
Lemma B.3.
@@
Suppose the statement holds for matrices of width <n<n and we now have a matrix $\mathbf{M}$ of width $n$. Let $L\subset V$ be the leaves of $T$. Let $T^{\prime}$ be the tree $T$ with these leafs removed, and let $\mathbf{M}^{\prime}\in\{0,1\}^{m\times(V\setminus L)}$ be matrix $\mathbf{M}$ with all columns corresponding to $L$ removed. Then for any $y\in V\setminus L$
@@
where equation 1 uses the fact that the path from root $r$ to $y$ is a path from root $r$ to $z$ with one extra edge $(z,y)$. Equation 2 uses the fact the path from $r$ to $z$ in $T$ also exists in $T^{\prime}$, and finally equation 3 follows by the induction hypothesis. Together, they prove the lemma. ∎
@@
We compute $\mathbf{N}^{(T)}v$ via Lemma˜B.2 which requires us to compute the vector $\vec{\sigma}\in\mathbb{R}^{E}$ with $\vec{\sigma}_{(y,c)}=\sum_{z\in V(T_{c})}v_{z}$ for all $(y,c)\in E$ where $T_{c}$ is the subtree of $T$ rooted at $c$. This can be done in $O(n)$ time via a simple “dynamic programming on trees” algorithm: For any $c\in V$, let $C$ be its children.
Then,
@@
Correctness. Without loss of generality, assume that for root $r$ of tree $T$ we have $\mathbf{M}_{r}=\vec{0}$. Otherwise, append a $\vec{0}$-column to $\mathbf{M}$, and add edge $(n+1,r)$ to $T$ with label $\Delta_{n+1,r}=\mathbf{M}_{r}-\mathbf{M}_{n+1}=\mathbf{M}_{r}$, and make $n+1$ the new root of $T$. This increases the tree weight by at most $m$, resulting in an additive $O(m)$ in the time complexity.
Append another entry to $v$ so that the dimensions match and $\mathbf{M}v$ is well-defined. By Lemma˜B.3, we thus have $\mathbf{N}^{(T)}v=\mathbf{M}v$.∎
B.2 Existence of Low Weight $\Delta$-labeled Minimum Spanning Trees
We first provide a proof of Lemma˜3.11 (Lemma 4.1 of Welzl [1988]), restated below for completeness. See 3.11 For this, we introduce the notion of $\epsilon$-nets and a crucial proposition of Welzl [1988] and Haussler and Welzl [1986].
Definition B.4 ($\epsilon$-net).
Let $(X,R)$ be a range space and $0\leq\epsilon\leq 1$. Let $A$ be a finite multiset of elements in $X$. Then, $N\subseteq A$ is an $\epsilon$-net of $A$ for $R$ if $\frac{|A\cap r|}{|A|}>\epsilon$ implies $r\cap N\neq\emptyset$ for all $r\in R$.
Proposition B.5 ([Welzl, 1988, Proposition 2.4]).
Let $(X,R)$ be a range space of VC-dimension $d\geq 1$ and let $0<\epsilon\leq 1$ be a real number. For every finite multiset $A$ of elements in $X$, there is an $\epsilon$-net $N$ of $A$ for $R$ such that $|N|\leq\lceil\frac{8d}{\epsilon}\log\frac{8d}{\epsilon}\rceil$.
Lemma B.6 (Lemma 2.2 of Welzl [1988]).
[…]
We are now ready to prove Lemma˜3.11, which we adapt from Welzl [1988].
Proof of Lemma˜3.11.
By the pigeonhole principle, it is sufficient to show the existence of a subset $N$ of $Q$ such that (i) $\pi^{*}(|N|)<n$ and (ii) if a pair of elements $\{x,y\}$ is crossed by more than $dmn^{-1/d}\log n$ ranges in $Q$, then $\{x,y\}$ is crossed by a range in $N$.
For $x\in X$, let $R_{x}=\{r\in R:x\in r\}$ and for $x,y\in X$ with $x\neq y$, let $R_{xy}$ be the set of ranges in $R$ that cross $\{x,y\}$. Since $(R,\{R_{x}|d\in X\})$ is the dual of $(X,R)$ it has VC-dimension at most $d$ by assumption. So by Lemma˜B.6, $(R,\{R_{xy}|x,y\in X,x\neq y\})$ has VC-dimension $O(d\log d)$.
Therefore, by Proposition˜B.5, for every $\epsilon>0$ and every multiset $Q$ of ranges in $R$, there is an $\epsilon$-net of $Q$ for $\{R_{xy}|x,y\in X,x\neq y\}$ with $|N|\leq\lceil\frac{O(d\log d)}{\epsilon}\log\frac{O(d\log d)}{\epsilon}\rceil$. Now, let $\epsilon=cdn^{-1/d}\log n$, where $c$ is some constant. Then, there is an $\epsilon$-net of $Q$ of size at most
$O(d\log d)\cdot\frac{n^{1/d}}{cd\log n}\log\left(O(d\log d)\frac{n^{1/d}}{cd\log n}\right)\leq c_{1}n^{1/d}$
where $c_{1}\leq O(1)$. For large enough $c$, we can get small enough $c_{1}$ such that $\pi^{*}(\lfloor c_{1}n^{1/d}\rfloor)<n$. For some constant $c$, we thereby get a $(cdn^{-1/d}\log n)$-net of $Q$ for the ranges $R_{xy}$ with $\pi^{*}(|N|)<n$. This implies that there is a pair of elements $\{x,y\}$ crossed by more than $cd|Q|n^{-1/d}\log n=cdmn^{-1/d}\log n$ ranges in $Q$. Finally, applying $d\leq\log n$ concludes the argument. ∎
See 3.14
Proof.
The crossing number of $R_{i}$ in T{T} counts the number of edges in T{T} such that one endpoint of the edge is in $R_{i}$ and the other endpoint is in $n]\setminus R_{i}$. Suppose the spanning tree $T$ on $X$ has a crossing number of $\kappa_{i}$ for just $R_{i}$ and let $w(T)=\sum_{i=1}^{m}\kappa_{i}$ be the weight of $T$ for $(X,R)$.
Perform a DFS traversal on T{T} and write the in-order traversal $\mathcal{I}$. We use the permutation $\pi(i)=\mathcal{I}_{i}$ for $i\in[n]$. Let $T^{\prime}$ be the corresponding linear tree. In the worst case, this vertex path crosses $R_{i}$ $2\kappa_{i}$ times, that is because DFS uses each edge $(x,y)$ of $T$ twice: once when visiting $y$ (going down one level in the tree) and once when it is done processing $y$ (DFS pops the current level from the stack).
Thus the total number of crossings $w(T^{\prime})$ for the linear tree $T^{\prime}$ is $w(T^{\prime})\leq\sum_{i=1}^{m}2\kappa_{i}\leq 2\kappa(T).$
B.3 Dynamic OMv Data Structure
In this section we give a black-box reduction from dynamic OMv to static OMv. In the static version, we are given a matrix $\mathbf{M}$ to preprocess. This matrix stays fixed, while we receive an online sequence of vectors and must repeatedly compute $\mathbf{M}v$. In the dynamic version, we can inform the data structure about changes to $\mathbf{M}$. Specifically, we are allowed to add/remove/change rows/columns of the matrix.
Using the reduction from dynamic to static, we obtain Theorem˜1.4.
See 1.4
@@ B.3 Dynamic OMv Data Structure
When there have been $2^{i}/2$ postponed deletions to any $\mathbf{M}^{(i)}$, then the columns are actually deleted and the respective static OMv data structure is reinitialized on the new $\mathbf{M}^{(i)}$. The remaining columns in $\mathbf{M}^{(i)}$ all exist in the current dynamic input $\mathbf{M}$, so again, the static OMv data structure receives a submatrix of $\mathbf{M}$ as input. For any one $i$, this takes amortized $O(P(n,2^{i})/2^{i})$ time which is subsumed by the complexity of handling insertions.
Observe that internally, our matrices $\mathbf{M}^{(i)}$ are always at most twice their intended size because of the yet to be removed columns. Thus the query complexity increases by at most a constant factor. ∎
We now prove Lemma˜B.7 (Row & Column Updates)
Proof.
Here, we extend Lemma˜B.8 to support row updates as well. The reduction follows the same idea as Lemma˜B.8.
Let $\mathbf{N}$ be the dynamic input matrix. We run $\log n$ copies of the column update data structure Lemma˜B.8 on matrices $\mathbf{N}^{(0)},...,\mathbf{N}^{(\log n)}$, with the $i$’th matrix having at most $2^{i}$ rows. When inserting new rows to $\mathbf{N}$, they are passed to $\mathbf{N}^{(0)}$. When $\mathbf{N}^{(i)}$ has more than $2^{i}$ rows, all rows are appended to $\mathbf{N}^{(i+1)}$. When rows are appended to any $\mathbf{N}^{(i)}$, the respective data structure is reinitialized. To handle deletions, we remove the respective row from the output until $2^{i}/2$ rows have been deleted from $\mathbf{N}^{(i)}$, then the rows are removed from $\mathbf{N}^{(i)}$ and the respective data structure is reinitialized. The amortized row update time thus becomes
@@
$O\left(\sum_{j=0}^{\log m}\sum_{i=0}^{\log n}\frac{P(2^{j},2^{i})}{2^{i}}\right)=O\left(\sum_{i=0}^{\log n}\frac{P(m,2^{i})}{2^{i}}\right)$
@@
which proves the lemma. ∎
We now prove Theorem˜1.4 (Dynamic OMv for corrupted VC-dimension $d$).
Proof.
The initialization cost of static OMv is $P(m,n)=\widetilde{O}(mn)$, and the query complexity is $Q(m,n)=\widetilde{O}(nm^{1-1/d}+m)$ by Theorem˜1.3. This complexity also holds for submatrices, i.e., $Q(a,b)=\widetilde{O}(nm^{2-1/d}+m)$ for any $a\times b$ sized submatrix.
@@
$O\left(\sum_{i=0}^{\log n}\frac{P(m,2^{i})}{2^{i}}\right)=\widetilde{O}\left(\sum_{i=0}^{\log n}\frac{m2^{i}}{2^{i}}\right)=\widetilde{O}(m)$
@@
$O\left(\sum_{j=0}^{\log m}\frac{P(2^{j},n)}{2^{j}}\right)=\widetilde{O}\left(\sum_{j=0}^{\log m}\frac{2^{j}n}{2^{j}}\right)=\widetilde{O}(n)$
@@
$\widetilde{O}(Q(m,n))\leq\widetilde{O}(nm^{1-\frac{1}{d}}+m),$
which proves the theorem. ∎
B.4 Extension to Structured Non-Boolean Matrices
In this section, we prove a result for a subquadratic-time accelerated matrix-vector multiplication for non-Boolean matrices, when the analogous Pollard pseudodimension is bounded. We also list further applications of this result.
Theorem B.9.
Suppose that the Pollard pseudodimension of the matrix $\mathbf{M}\in\mathbb{R}^{m\times n}$ is $d$, and that the number of distinct values per column is $\leq T$. Then, after an $\widetilde{O}(Tmn)$ time preprocessing, there is a data structure that, upon receiving a vector $v\in\mathbb{R}^{n}$, returns $\mathbf{M}v$ in time $O(Tnm^{1-1/d}+m)$.
Remark B.10.
When $\mathbf{M}$ is the weight matrix of a neural network and $v$ is the activation output from the previous layer of the neural network, the matrix-vector product $\mathbf{M}v$ corresponds to running inference on $v$ with the neural network $\mathbf{M}$. This is a key step in a number of applications, including online optimization in unknown no-linear systems [Lin et al., 2023a] and inference in large language models. Dehghankar et al. [2024] studies the latter application with ternary matrices $\mathbf{M}\in\{-1,0,1\}^{n\times m}$ and shows that they are amenable to $O(n^{2}/\log n)$ multiplication. If $\mathbf{M}$ has bounded Pollard pseudodimension $d$, then as $T=3$, we obtain a polynomial acceleration on this result: by Theorem˜B.9, the matrix-vector product can be computed in $\widetilde{O}(mn^{1-1/d}+n)$ time.